OpenStudy (anonymous):

How do you simplify the following, can you please show me a step by step process? Thank you

OpenStudy (anonymous):

\[(-\sqrt{3}-1)/(1+(-\sqrt{3})(1))\]

OpenStudy (mertsj):

\[\frac{-\sqrt{3}-1}{1-\sqrt{3}}\times\frac{1+\sqrt{3}}{1+\sqrt{3}}=\frac{-\sqrt{3}-3-1-\sqrt{3}}{1-3}\]

OpenStudy (anonymous):

\[2+\sqrt{3}\]

OpenStudy (anonymous):

- (√3+1) * √3+1) / (1 -√3) (√3+1) = - √3+1) ^2/ -2

OpenStudy (mertsj):

\[\frac{-4-2\sqrt{3}}{-2}=2+\sqrt{3}\]